### 24 Jan 2024 v1

# Scalable Link Prediction on Large-Scale Heterogeneous Graphs with Large Language Models

Abstract
Exploring the application of large-scale language models to graph learning is a novel endeavor. However, the vast amount of information inherent in large graphs poses significant challenges to this process. This paper focuses on the link prediction task and introduces LPNL (Link Prediction via Natural Language), a framework based on a large language model designed for scalable link prediction on large-scale heterogeneous graphs. We design novel prompts for link prediction that articulate graph details in natural language. We propose a two-stage sampling pipeline to extract crucial information from large-scale heterogeneous graphs, and a divide-and-conquer strategy to control the input token count within predefined limits, addressing the challenge of overwhelming information. We fine-tune a T5 model based on our self-supervised learning designed for for link prediction. Extensive experiments on a large public heterogeneous graphs demonstrate that LPNL outperforms various advanced baselines, highlighting its remarkable performance in link prediction tasks on large-scale graphs.
1 Introduction
The emergence of large language models (LLMs) has elevated the field of natural language processing to unprecedented levels Devlin et al. (2018); Raffel et al. (2020); Brown et al. (2020). Recently, some research Fatemi et al. (2023); Ye et al. (2023) has explored the use of LLMs to address the learning on graphs. A popular paradigm of using LLMs on graph learning is to transform graph problems and structures into texts, and then feed the texts to LLMs to obtain the predictions. Figure 1 shows two pipelines for link prediction Daud et al. (2020); Lichtenwalter et al. (2010) based on Graph Neural Networks (GNNs) and LLMs. The former employs GNNs for graph learning, followed by a matcher to compute similarity to link predict Zhang and Chen (2018); Cai et al. (2021). Meanwhile, the latter generates textual prompts based on sampled subgraphs and utilizes LLMs for prediction.
In contrast to the conventional approach of GNNs that aggregate encoded features via message passing Hamilton et al. (2017); Abu-El-Haija et al. (2019), LLMs exhibit superior learning capabilities and enhanced flexibility Tamkin et al. (2021); Liang et al. (2022). However, it remains under explored on how to effectively learn from larger and more complex graphs with LLMs, which poses high challenges to capturing distant information and richer semantics. An example is the link prediction on large-scale graphs: as the number of candidate nodes or description nodes for prediction grows, so does the text fed into LLMs. In this case, extensive inputs become unfeasible due to token length limitations.
Figure 1: GNNs and LLMs based pipelines for link prediction
In this work, we explore the scalable link prediction with large language models on large-scale heterogeneous graphs(LSHGs). LSHGs Shi et al. (2016); Wang et al. (2022) can be considered as networks of connections that resemble the real world more closely than conventional graph structures, such as social network Robins et al. (2007) or academic networks Kong et al. (2019). The key challenges can be succinctly described as follows: 1) how to fomulate the prompt template for scalable link prediction task. 2) how to find out crucial information on LSHGs by sampling, enabling LLMs to capture it within limited inputs. 3) how to address lengthy prompts generated by an excess of candidate nodes.
To tackle the above challenges, we propose LPNL (Link Prediction via Natural Language), a large language model based framework for scalable link prediction on large-scale graphs. We design novel prompts for link prediction that articulating graph details in natural language. This involves establishing a selective query prompt template, furnishing a description of the link prediction task, and integrating heterogeneous information concerning the source and candidate nodes.
In dealing with vast amounts of relevant graph information within LSHGs, LPNL’s focus lies on the crucial segments. This approach helps us avoid the interference of superfluous contextual information while ensuring compliance with specified token limitations. We design a two-stage sampling pipeline that utilizes normalized degree-based heterogeneous subgraph sampling and personalized pagerank-based ranking. This pipeline selects more crucial node information from the graph, ensuring that the model focuses more on more informative nodes.
Figure 2: The framework of LPNL on heterogeneous graphs. For an input heterogeneous graph with link prediction tasks, perform two-stage sampling on the source node and each candidate node, resulting in the generation of corresponding anchor nodes. Generate prompts based on anchor nodes, input them into a large language model, reconstruct candidates based on the prediction results, create new prompts for another round of prediction, and repeat this divide-and-conquer process to obtain a unique link prediction result.
With a large number of candidate nodes and only a single link prediction query, the token length constraints make it challenging to fully describe all candidate node information. For this situation, we utilize a divide-and-conquer method. We partition the original node set into multiple equally sized smaller sets, sequentially input them into the link prediction pipeline to obtain partial answers, and recursively reconstruct the candidate set to obtain the final answer.
We conduct extensive experiments on the Open Academic Graph (OAG) Huang et al. (2020) and fine-tune the language models T5 Raffel et al. (2020) to serve as the backbone models on the OAG for LPNL. The results demonstrate that LPNL outperforms various enhanced GNNs-based baselines significantly in performance. Furthermore, through extensive experimentation, we discover that LPNL also exhibits remarkable few-shot capability. Unlike traditional models, LPNL’s training merely requires simple alignment formatting, enabling swift convergence in predictions and showcasing superior performance compared to conventional models. Additionally, experiments demonstrate the model’s robust knowledge transferability, maintaining consistent performance across various cross-domain tasks. This further underscores the immense potential of large language models in addressing graph learning.
2 The LPNL Architecture
In this section, we introduce the details of our proposed Link Prediction via Natural Language, i.e. LPNL, a framework utilizing natural language to solve link prediction task on large-scale heterogeneous graphs. We start with the notation setup, followed by the prompt design, the sampling methods, and our divide-and-conquer and self-supervised strategy with more details.
Figure 3: An example of heterogeneous graph
2.1 Preliminary
Figure 3 showcases heterogeneous type nodes and their heterogeneous relationships sourced from OAG. Formally, a heterogeneous graph is denoted by $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{A},\mathcal{R}\}$, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of nodes and edges (links), respectively. Each node $v\in\mathcal{V}$ and each link $e\in\mathcal{E}$ are associated with their mapping function $\phi(v):v\rightarrow\mathcal{A}$ and $\varphi(e):e\rightarrow\mathcal{R}$. $\mathcal{A}$ and $\mathcal{R}$ represent the sets of nodes and edge types.
Given a source node $s$ and a set of candidate nodes $\mathcal{C}=\{c_{1},c_{2},...,c_{n}\}$ satisfying a existed meta-relation $\langle\phi(s),\varphi(e),\phi(c_{i})\rangle$ where $e\in\mathcal{E}$ and $c_{i}\in\mathcal{C}$. Then a standard link prediction task on heterogeneous graphs is to predict a candidate node $c\in\mathcal{C}$ for a source node $s$ with the highest probability of $\langle s,e,c\rangle$.
Finally, let $\mathcal{G}_{sub}^{h}(v)=\{\mathcal{V}_{v}^{h},\mathcal{E}_{v}^{h},\mathcal{A}_{v}^{h},\mathcal{R}_{v}^{h}\}$ denote the $h$-hop ego-subgraph around $v$, consisting of $h$-hop neighbor nodes of $v$ and all interconnecting edges. We also denote $\mathcal{N}^{h}(v)$ as the set of all neighbor nodes on $\mathcal{G}_{sub}^{h}(v)$, which means $\mathcal{N}^{h}(v)=\{v^{\prime}|v^{\prime}\in\mathcal{V}_{v}^{h},v^{\prime}\neq v\}$. And $\mathcal{AN}^{h}_{k}(v)$ is denoted as the sequence of top-$k$ anchor nodes which is selected from $\mathcal{N}^{h}(v)$. Note that the above definitions are all heterogeneous.
2.2 Prompt Design for Link Prediction
In order to comprehensively represent the link prediction task along with the essential graph information, we meticulously design a uniform prompt template $\mathcal{T}(\cdot)$ for heterogeneous link prediction. Its fundamental mode involves a selective query, presenting both the link prediction problem description and information regarding the source node and candidate nodes. It prompts the large language models to identify the node most likely to be linked within the candidate set.
First, we define $d(v)$ as the description of node $v$, which consists of a sequence of textual features of itself and also its top-$k$ anchor nodes:
$d(v)=\{v:\mathcal{S}_{v}\}\text{ is related with }\sum_{i=1}^{k}\{v^{\prime}_{i}:\mathcal{S}_{v^{\prime}_{i}}\}$ (1)
where $\mathcal{S}_{v}$ denotes the textual description of node $v$ and $v^{\prime}_{i}$ represents the anchor node of node $v$ satisfying $v^{\prime}_{i}\in\mathcal{AN}^{h}_{k}(v)$.
Subsequently, given a source node $s$ and the set of candidate nodes $\mathcal{C}$, we formally obtain the link prediction prompt template as follows:
$\mathcal{T}(s,\mathcal{R},\mathcal{C})=q(\mathcal{R})+d(s)+\sum_{i=1}^{n}d(c_{i}|c_{i}\in\mathcal{C})$ (2)
where $\mathcal{R}$ is the relation type between the source node and candidate nodes and $n$ is the number of candidate nodes. And $q(\mathcal{R})$ represents a link prediction query, e.g., "which $\phi(c)$ is linked by $\phi(u)$?". Notably, in the above equation, the addition operators are redefined as the textual concatenation with separators. To better enable the large-scale model in distinguishing between various types of heterogeneous nodes, we additionally assign distinct type identifiers to the backend of each node, e.g. a paper node: <p>[PA]. We also show a more intuitive prompt template example in Figure 1.
Our designed prompts do not encompass the link information in the graph. Instead, we describe the key nodes in textual form based on their order of importance. This is because of the complexity of inter-node connections, which often result in redundant contexts Fatemi et al. (2023), making it challenging for the LLMs to comprehend. Consequently, LLMs might reducing the emphasis on the node features, which are pivotal for our tasks. Nonetheless, the links among heterogeneous nodes remain crucial as they reflect their relationships and node significance. To address this, we utilize the following two-stage sampling approach to prioritize critical nodes, thereby better encapsulating graph information.
2.3 Two-Stage Sampling
In the previous subsection, we designed the unified prompt template for link prediction. However, as graph data becomes more complex, resembling the real world, employing a single prompt engineering approach becomes challenging in addressing practical application problems.
Firstly, in large-scale heterogeneous graphs, if Eq.(1) attempts to represent the node $v$ information using its all $h$-hop neighbors, i.e. $k=\lvert\mathcal{V}_{v}^{h}\rvert$ formally, it will lead to an uncontrollable prompt length. Anothor issue arises due to substantial variations in the degrees of different node types. For example, the number of nodes in the $h$-hop subgraph around a paper-type node $\lvert\mathcal{N}^{h}($<p>)|$\rvert$) | is significantly smaller than that around a field-type node $\lvert\mathcal{N}^{h}($<f>)|$\rvert$) |. The two problems pose significant challenges to the input and contextual comprehension of large models.
In this paper, we provide a two-stage sampling pipeline. The first stage aims to sample subgraphs $\mathcal{G}_{sub}^{h}(v)$ based on normalized degree from large-scale heterogeneous graphs while mitigating sampling bias caused by heterogeneous types. And then we obtain the top-$k$ anchor nodes sequence $\mathcal{AN}^{h}_{k}(v)$ through the second stage sampling with personalized pagerank to generate $d(v)$ in Eq.(1).The further details are as follows.
Normalized Degree based Sampling Inspired by previous studies  Hu et al. (2020); Leskovec and Faloutsos (2006), we sample heterogeneous subgraphs based on normalized degree. Specifically, it specifies the sampling probability of each hop’s neighbors as their normalized degree. Normalized degree is defined as the current node’s degree normalized among all nodes of the same type in the same layer awaiting sampling. Therefore, for the $l$-th layer subgraph sampling around central node $s$, the sampling probability of node $v$ can be described as follows:
$prob^{l}_{s}(v)=\frac{deg(v)^{2}}{\sum{deg(u)^{2}}}$ (3)
where node $u$ represents the neighbor node at the $l$-th layer within the subgraph, satisfying $u\in\mathcal{V}_{s}^{h}-\mathcal{V}_{s}^{h-1}$ and $\phi(v)=\phi(u)$.
The normalized degree based sampling ensures that our first stage of sampling do not ignore the differences between node types so that it does not lead to bias against certain node types, e.g., nodes with higher degrees were not indiscriminately considered more important. Thus, a similar number of different types of nodes in the subgraphs was maintained, preserving richer semantic information. Furthermore, previous studies have demonstrated that leveraging up to 3-hop connectivity is effective for achieving excellent performance Kipf and Welling (2016); Veličković et al. (2017); Hamilton et al. (2017). However, extending the information beyond 3-hop generally has a marginal impact on improvement and, in some cases, may even result in negative effects Cai and Wang (2020); Zhang et al. (2021). Thus, we set the maximum value for multi-hop to 2-hop or 3-hop to implement the two-stage sampling approach.
Sampling with Personalized PageRank Through the sampling in the first stage, the heterogeneous subgraphs we obtain eliminate biases between different types, allowing all types of nodes to be compared regarding their importance on an equal footing. In the second stage, we directly compute the importance of all neighbor nodes within the subgraph $\mathcal{G}_{sub}^{h}(v)$ for the source node $s$ using Personalized PageRank(PPR) Bojchevski et al. (2020); Vattani et al. (2011). We then obtain the PPR vector $\vec{\pi_{s}}$ for the source node $s$ by iteratively updating the following:
$\vec{\pi_{s}}=\alpha*\vec{e_{s}}+(1-\alpha)*A^{\top}D^{-1}\vec{\pi_{s}}$ (4)
where $\alpha$ denotes the damping factor, $A$ stands for the adjacency matrix, $D^{-1}$ denotes the diagonal degree matrix and $\vec{\pi_{s}}$ signifies the unit vector.
This work employs a queue-based implementation of the equivalent random walk Spitzer (2013); Wu et al. (2021) to approximate PPR. Subsequently, the top-$k$ anchor nodes sequence $\mathcal{AN}^{h}_{k}(s)$ is obtained based on the ranking derived from PPR, which characterizes the top-$k$ neighbor nodes that are most critical for the source node $s$ within the whole hetero-graph.
The two-stage sampling restricts the generated link prediction prompt length to suit large model inputs while maximizing the retention of crucial neighborhood information pertaining to the target node within the subgraphs. It also makes use of the structural information on the graph, so that the generated anchor nodes $\mathcal{AN}^{h}_{k}(s)$ can be seen as a hub converting the graph structure into textual descriptions. This enables our prompts generated in Sec. 2.2 to encompass not only node features but also implicit structural information.
Dataset ##\##nodes ##\##edges ##\##papers ##\##authors ##\##fields ##\##venues ##\##institutes ##\##P-A ##\##P-F ##\##P-V ##\##A-I ##\##P-P
CS 11,918,983 107,263,811 5,597,605 5,985,759 119,537 27,433 16,931 15,571,614 47,462,559 5,597,606 7,190,480 31,441,552
Mater 4,552,941 42,161,581 2,442,235 2,005,362 79,305 15,141 10,898 5,582,765 19,119,947 2,442,235 2,005,362 13,011,272
Engin 5,191,920 36,146,719 3,239,504 1,819,100 99,444 19,867 14,005 3,741,135 22,498,822 3,239,504 1,819,100 4,848,158
Chem 12,158,967 159,537,437 7,193,321 4,748,812 183,782 19,142 13,910 16,414,176 57,162,528 7,193,321 4,748,812 74,018,600
Table 1: Open Academic Graph (OAG) Statistics.
2.4 Divide-and-Conquer Prediction
While the previous subsection’s sampling pipeline addressed the potential issue of prompts length caused by Eq.(1), a careful observation of Eq.(2) reveals that an excessive number of candidate nodes in link prediction, denoted as $|\mathcal{C}|$, also makes the prompt length uncontrollable. Especially in large-scale heterogeneous graphs, the number of candidate nodes is high, making it difficult to describe all candidate nodes within a window-sized prompt. For instance, in a link prediction task with 100 candidate nodes, each node requires an average of approximately 250 tokens in the prompt for description. This would amount to a total of 25,000 tokens needed to describe all candidate nodes. This far exceeds the maximum token limit for a usual LLM’s context window. Furthermore, the excessive number of candidate nodes leads to redundant contexts, making it challenging for the LLMs to comprehend the input text.
LPNL avoids the aforementioned token overload by employing a divide-and-conquer strategy Bentley (1980); Smith (1987). Firstly, we set a candidate length limit $\mathcal{L}$, ensuring that the length of all processed candidate sets does not exceed $\mathcal{L}$. Specifically, for an original candidate set $\mathcal{C}_{ori}$ of length $|\mathcal{C}_{ori}|$ where $|\mathcal{C}_{ori}|>\mathcal{L}$, we randomly divide it into $\lceil\frac{\mathcal{C}_{ori}}{\mathcal{L}}\rceil$ sub-candidate sets, i.e., $\mathcal{C}_{1}^{1},\mathcal{C}_{2}^{1},...,\mathcal{C}_{\lceil\frac{\mathcal{C}_{ori}}{\mathcal{L}}\rceil}^{1}$, satisfying $max(|\mathcal{C}_{1}^{1}|,|\mathcal{C}_{2}^{1}|,...,|\mathcal{C}_{\lceil\frac{\mathcal{C}_{ori}}{\mathcal{L}}\rceil}^{1}|)\leq\mathcal{L}$. As illustrated in Figure 1, by employing the fine-tuned large language model to predict the linked node with the maximum link probability for each sub-candidate set, we can subsequently eliminate low-probability candidate nodes. This process leads to more refined candidate sets by candidates reconstruction. And following this divide-and-conquer process by redividing candidate sets and making predictions, ultimately, we can obtain a unique prediction answer for the entire original candidate set $\mathcal{C}_{ori}$. Figure 4 provides a more intuitive illustration of the divide-and-conquer prediction, allowing us to observe the descent of candidate nodes and prompt tokens throughout the process.
Figure 4: Divide-and-conquer Prediction
2.5 Self-supervised Learning based Finetune
As a more relevant graph structure, large-scale heterogeneous graphs lack labelled data. LPNL supports self-supervised learning based finetune for large language models without relying on labels. During the end-to-end prompt tuning, it automatically constructs a candidate set containing ground truth, aligned with downstream prediction formats. And the ground truth is used as the correct answer for link prediction. To ensure training correctness, the ground truth appears randomly within the candidate sequence, and notably, during the heterogeneous subgraph sampling process in Section 2.3, the edges between the ground truth and the source node are masked.
3 Experiments
3.1 Experiment Settings
Dataset Metric GraphSage HGT RGCN GCN GAT LPNL $\Delta$
CS NDCG .814$\pm$.025 .847$\pm$.042 .843$\pm$.056 .887$\pm$.031 .911$\pm$.033 .985$\pm$.008 $\uparrow 8.12\%$
MRR .640$\pm$.045 .712$\pm$.024 .685$\pm$.056 .727$\pm$.032 .797$\pm$.051 .939$\pm$.018 $\uparrow 17.81\%$
Hits@1 .469$\pm$.012 .562$\pm$.022 .532$\pm$.056 .568$\pm$.011 .686$\pm$.014 .894$\pm$.004 $\uparrow 30.32\%$
Mater NDCG .765$\pm$.017 .841$\pm$.034 .854$\pm$.042 .818$\pm$.016 .897$\pm$.065 .954$\pm$.014 $\uparrow 6.35\%$
MRR .519$\pm$.052 .643$\pm$.031 .665$\pm$.027 .667$\pm$.036 .747$\pm$.058 .881$\pm$.011 $\uparrow 17.93\%$
Hits@1 .278$\pm$.016 .447$\pm$.019 .476$\pm$.028 .524$\pm$.032 .597$\pm$.018 .809$\pm$.007 $\uparrow 35.51\%$
Engin NDCG .798$\pm$.021 .876$\pm$.022 .874$\pm$.061 .912$\pm$.041 .913$\pm$.037 .977$\pm$.017 $\uparrow 7.01\%$
MRR .570$\pm$.027 .691$\pm$.041 .699$\pm$.034 .747$\pm$.023 .769$\pm$.041 .917$\pm$.017 $\uparrow 16.14\%$
Hits@1 .342$\pm$.023 .506$\pm$.018 .523$\pm$.056 .583$\pm$.021 .624$\pm$.011 .858$\pm$.012 $\uparrow 37.50\%$
Chem NDCG .821$\pm$.015 .863$\pm$.015 .835$\pm$.036 .893$\pm$.017 .899$\pm$.023 .955$\pm$.018 $\uparrow 6.23\%$
MRR .649$\pm$.034 .724$\pm$.027 .678$\pm$.031 .749$\pm$.023 .780$\pm$.029 .872$\pm$.038 $\uparrow 11.79\%$
Hits@1 .485$\pm$.024 .523$\pm$.031 .530$\pm$.016 .609$\pm$.020 .667$\pm$.022 .792$\pm$.007 $\uparrow 18.74\%$
Table 2: Experimental results of different methods over the four datasets.
Models We finetune Flan-T5 Chung et al. (2022) as the backbone language model for our LPNL. The numbers of sampling hops $h=3$, top anchor nodes sequence $k=50$ and candidate length limit $\mathcal{L}=3$, mentioned in Section 2.3 and 2.4 is used for all LPNL.
Datasets We conducted all experiments on the OAG dataset Huang et al. (2020), known as one of the largest publicly available heterogeneous graphs, comprising 178 million nodes and 2.236 billion edges. It includes five types of nodes (denoted as papers (P), authors (A), venues (V), institutes (I) and fields (F)) and their interrelations. In our specific experiments, we utilized four representative domain-specific subgraphs from OAG: Computer Science (CS), Material Science (Mater), Engineering (Engin) and Chemistry (Chem) Jiang et al. (2021). The graph statistics are listed in Table 1. We partition each dataset into finetune, validation, and test sets based on distinct time periods. Specifically, in the OAG dataset, papers are published between 1900 and 2019. Consequently, we utilize publications preceding 2015 for finetune, data from 2015 to 2016 for validation, and information from 2016 onwards for testing.
Task We consider the real-world link prediction tasks to evaluate the performence of our LPNL: author name disambiguation Ferreira et al. (2012), which is a fundamental challenge for curating academic publication and author information, as duplicated names widely exist in our lives. Specifically, given a paper, the objective is to predict the true author who has a genuine link with that paper among all authors with the same name.
Baselines We select a series of supervised baselines, all of which are advanced graph neural network models. These include GCN Li et al. (2018), GraphSage Hamilton et al. (2017) and GAT Veličković et al. (2017), designed for homogeneous graphs, as well as RGCN Schlichtkrull et al. (2018) and HGT Hu et al. (2020), tailored for heterogeneous graphs.
3.2 Overall Performance
In this experiment, we compare the T5 model as the backbone version of LPNL to advanced GNN based baseline models across the four domain-specific subgraphs. We finetune the model separately across various subgraphs and tasks and evaluate the performance of the models in link prediction. The experimental results of the proposed method and baselines are summarized in Table 2. All experiments for the author name disambiguation task over all datasets are evaluated in terms of NDCG, MRR and Hits@1 Li (2022); Liu et al. (2009).
The results show that in terms of all three metrics, the proposed LPNL significantly and consistently outperforms all baselines for all tasks on all datasets. Overall, our LPNL consistently yields the best performance among all methods, leading to an average improvement of 6.93%, 15.92% and 30.52%, compared to the second best baseline method. Surprisingly, LPNL exhibited significant improvements in all settings, particularly in the Hit@1 metric. The substantial leap in achieving correct predictions with just a single attempt holds significant implications for practical applications. These improvements over GNNs indicate the efficacy of our proposed LPNL in enabling large language models to comprehend link prediction tasks within complex graphs and large language models have tremendous potential in addressing graph-related problems.
Figure 5: Cross-Domain Transfer Results
3.3 Cross-Domain Knowledge Transfer
To explore the generalization capabilities of LPNL, we set up experiments for cross-domain knowledge transfer. Specifically, we fine-tuned the T5 model using LPNL on a subgraph corresponding to one domain and subsequently conducted testing on subgraphs from other domains. The experimental outcomes, visualized in Figure 5 as a heatmap, reveal that in most instances, the model exhibits optimal performance when fine-tuned within its original domain. Surprisingly, models fine-tuned on other domains also demonstrate remarkably strong performance, often closely matching or even surpassing the best performance achieved by fine-tuning within the original domain (e.g., Mater-Engin). This highlights the robust knowledge transferability of our approach, showcasing minimal sensitivity to domain-specific information and demonstrating robust generalization capabilities.
Figure 6: Comparison of few-shot results
3.4 Few shot cability
The extensive pretraining of large language models across various natural language tasks has endowed them with robust reasoning and generalization capabilities. In contrast to traditional GNN models, they require minimal training samples to converge and exhibit superior performance. We further investigates the few-shot capabilities of LPNL by comparing it with the top-performing homogeneous GNN and heterogeneous GNN in terms of overall performance. Configure the evaluate results to be printed every 1 batch, with each batch consisting of 50 link prediction tasks. Compare the few-shot results for the first 20 batches. The results in Table 6 demonstrate that our LPNL swiftly converges with minimal sample fine-tuning, displaying comparable performance to the best fine-tuning outcomes. This showcases the portability of large language models in addressing graph-related tasks.
Method NDCG MRR Hits@1
LPNL 97.86 94.37 89.47
w/o graph 68.98 50.51 35.97
w/o Stage 1 76.31 57.29 41.83
w/o Stage 2 87.67 70.67 53.33
Table 3: Ablation Study Results of Sampling Methods
3.5 Ablation Study
We conducte an ablation study on CS dataset to evaluate the effectiveness of our approach in employing large language models combined with graph knowledge strategies. We compare the performance among different versions of sampling methods: the standard LPNL, a version without any graph information, and another two sampling versions, each independently utilizing distinct stages. As illustrated in Table 3, the model’s performance significantly diminishes when graph information is excluded. Furthermore, the performance of the versions without stage 1 and stage 2 shows a notable gap compared to LPNL. It indicates that employing our designed two-stage sampling pipeline enables LPNL to capture crucial information within the graph after balanced heterogeneous sampling, resulting in improved predictive outcomes.
 Hop Info NDCG MRR Hits@1
Multi-hop 97.86 94.37 89.47
1-hop 94.15 91.56 85.09
Table 4: Ablation Study Results of Multi-hop
In our experiments, two critical operations contributing significantly to the outstanding performance of LPNL in link prediction are multi-hop and anchor nodes, which provide essential information to the LLMs. To assess the impact of these two key components on model performance, we conducted another ablation experiments, and the results are presented in Table 4 and 5. It shows that incorporating multi-hop and more anchor nodes information can both enhance the LPNL’s performance. However, further experiments indicate that increasing the number of multi-hop and anchor nodes beyond a certain threshold does not lead to significant performance improvement. On the contrary, it may result in additional costs without notable benefits.
 Top-k Info NDCG MRR Hits@1
Top-30 92.86 76.48 62.05
Top-50 97.86 94.37 89.47
Top-70 98.06 93.84 88.69
Table 5: Ablation Study Results of Anchor Nodes
4 Related Work
Graph Representation Learning Based on GNNs Graph Neural Networks (GNNs) are the forefront of graph representation learning methods and have gained significant popularity across a range of graph-related tasks Wu et al. (2020); Zhou et al. (2020). In these tasks, such as node classification and link prediction, GNNs-based approaches usually preprocess the corresponding text by a language model and encode the resulting embedding as node features. The final node representation is then obtained by aggregating the neighborhood features through spectral methods Bruna et al. (2013); Defferrard et al. (2016) and message passing Abu-El-Haija et al. (2019); Hamilton et al. (2017); Schlichtkrull et al. (2018). Besides, some studies have attempted to propose the GNNs architectures on heterogeneous graphs Dong et al. (2020); Wang et al. (2019); Hu et al. (2020). Notably, influenced by large language models, recent studies Sun et al. (2023); Huang et al. (2023)have explored the potential of GNNs in prompt learning. And there have also been attempts Ioannidis et al. (2022); Zhao et al. (2022) to explore collaborative training between Language Models and GNNs.
Large Language Models with Graph Knowledge The emergence of large language models (PLMs) has propelled natural language processing (NLP) to new heights. Qiu et al. (2020). For example, models like BERT Devlin et al. (2018) and T5 Raffel et al. (2020) demonstrate excellent performance in a wide range of downstream tasks, such as text classification and question answering. Besides, some works  Zhang et al. (2019); Liu et al. (2022, 2020)attempts to inject external graph knowledge into LLMs, thus enabling LLMs to gain the ability to solve problems on graphs Recently, due to the powerful inferential capabilities of large language models, a burgeoning body of work Fatemi et al. (2023); Ye et al. (2023); Liu et al. (2023) attempt to utilize natural language descriptions of graph features, employing generated prompts to instruct large language models in addressing various problems on graphs.
5 Discussion
From our experiments, we found that describing graphs using natural language does not follow the principle of ’more information is better.’ Sampling more nodes can introduce additional information, but it may lead to information redundancy, resulting in a decline in the inferential capabilities of large models. Therefore, the key lies in the setting of the sampling and divide-and-conquer length limits, which should align with the input window size and inferential capabilities of the large language model. While designing prompts, we observe that complex relationships between nodes are challenging to articulate in text, especially in large or dense graphs, potentially leading to redundant contexts. LPNL leverages structural information during the sampling phase and, in the prompt generation, only conveys information about the sampled nodes. This approach aims to minimize context redundancy while maximizing the utilization of graph information.
6 Conclusion
In this paper, we explore, for the first time, the application of large language models to address the link prediction task on large-scale heterogeneous graphs. We introduce LPNL, a large language model based framework for scalable link prediction on large-scale graphs. We design specific prompt templates for the link prediction task and generate the prompts based on anchor nodes obtained through a two-stage sampling approach. These prompts are then input to the large language model for predictions. To tackle the token overload issue arising from an excessive number of candidate nodes, we employ a divide-and-conquer strategy. Empirical evaluations demonstrate that LPNL achieves significant improvements compared to GNN baselines, showcasing its robust capability in cross-domain knowledge transfer and few-shot learning scenarios.